Even the best vacuum pumps cannot lower the pressure in a container below 10−15atm. How many molecules of air are left in each cubic centimeter in this "vacuum”? Assume that the temperature is 274 K.
Answer #1

answer) from the gas eqn we have the formula

PV=nRT

n=PV/RT

P=10-15atm

V=1cm3=0.001L

R=0.08205746Latm

T=274K

so plugging the values we have

n=10-15*0.001L/0.08205746*274=4.4477*10-20mol

now we have to multiply this avogadros number

n=4.4477*10-20mol*6.022*1023molecules/mol=2.68*104 moles

so the answer is 26784 molecules or 3.0*104 molecules or 2.7*104 molecules or 2.68*104 molecules.